determine the rate of reaction if the concentration of H2 and Cl2 are 2.5M and 3.0 M Respectively?
1
Expert's answer
2018-05-24T07:38:07-0400
.c (H2) =2.5 M
.c(Cl2) = 3.0 M
V reaction=?
The equation of the reaction between H2 and Cl2 is
H2+Cl2=2 HCl
The rate of the reaction according to the law of acting masses
V=k[H2] [Cl2] where k-reaction rate constant
V=k×2.5×3.0=7.5 k